Libby is an hourly employee who earns $14.32/hour. She is paid overtime only for hours worked past 40 in one week. During a weekly pay period, she worked the following hours:

Day Hours Minutes
Monday 8 47
Tuesday 8 22
Wednesday 6 25
Thursday 9 6
Friday 7 58

What is the difference between her pay using the quarter-hour and the hundredth-hour methods? Which method pays Libby the greater amount? (Do not round interim calculations. Round final answer to two decimal places.)
Answer #1

Solution a:

Total hours worked by Libby = 8 hours, 47 minute + 8 hours, 22 minutes + 6 hours, 25 minutes + 9 hours, 6 minutes + 7 hours, 58 minutes

= 40 hours, 38 minutes

Total time worked using quarter hour method = 40 hours + 45/60 = 40.75 hours (Rounded off to nearest quarter hour)

Pay using quarter hour method = 40*$14.32 + 0.75*$14.32*1.50 = $588.91

Total time worked using hundredth hour time = 40 hours + 38/60 = 40.63 hours

Pay using hundredth hour method = 40*$14.32 + 0.63*$14.32*1.50 = $586.33

difference between her pay using the quarter-hour and the hundredth-hour methods = $588.91 - $586.33 = $2.58

Quarter hour method pays libby the greater amount.